# Communication-Efficient Soft Actor-Critic Policy Collaboration via Regulated Segment Mixture in Internet of Vehicles

### v2

@@ I-D Paper Organization
The remainder of this paper is organized as follows. In Section II, we present preliminaries of SAC algorithm and main notations used in this paper. Then, we introduce the system model and formulate the problem in Section III. Afterwards, we provide the mixed performance improvement bound theorem of FRL communication under MERL in Section IV, and elaborate on the details of the proposed RSM-MASAC algorithm in Section V. In Section VI, we present the simulation settings and discuss the experimental results. Finally, Section VII concludes this paper.
@@ II Preliminary
$s_{t}^{(i)},a_{t}^{(i)},r^{(i)}_{t}$ Local state, individual action and reward of agent $i$ at time step $t$
@@
$p,P$ Index of segments, $p=1,2,\cdots,P$
$\Omega_{i}$ Set of one-hop neighbors within the communication range of agent $i$
$\zeta$ Mixture metric of current policy parameter vector and referential policy parameter vector
$\alpha$ Temperature parameter of policy entropy
$\varrho$ Target smoothing coefficient of target $Q$ networks
$\kappa$ Number of model replicas
$U$ Communication interval determined by specified iterations of the local policy
$\upsilon$ Transmission bits of the policy parameters
$\psi$   Communication consumption
@@
$\vspace{-0.2em}\max\eta(\pi)=\mathop{\mathbb{E}}_{s_{0}\sim\rho_{0}}\left[V^{\pi}(s_{0})\right]$ (1)
@@
    $V(s_{t})$ $=\mathbb{E}_{a_{t}\sim\pi}[Q(s_{t},a_{t})]+\alpha H(\pi(\cdot|s_{t}))$ (4)
@@
  • In the soft policy improvement step, the goal is to find a policy $\pi_{\text{new}}$ superior than the current policy $\pi_{\text{old}}$, in terms of maximizing (1). Specifically, for each state, SAC updates the policy as
@@
    $=$ $\mathop{\arg\max}_{\pi\in\Pi}\mathbb{E}_{a_{t}\sim\pi}\left[Q^{\pi_{\text{old}}}(s_{t},a_{t})-\alpha\log\pi(a_{t}|s_{t})\right]$
    where $\mathrm{D}_{\mathrm{KL}}(\cdot)$ denotes the KL divergence, while the partition function $Z^{\pi_{\text{old}}}$ normalizes the distribution. Since $Z^{\pi_{\text{old}}}$ has no contribution to gradient with respect to the new policy, it can thus be ignored. As unveiled in Appendix A, through the update rule of (5), $Q^{\pi_{\text{new}}}(s_{t},a_{t})\geq Q^{\pi_{\text{old}}}(s_{t},a_{t})$ is guaranteed for all $(s_{t},a_{t})\in\mathcal{S}\times\mathcal{A}$. Notably, the proof of soft policy improvement detailed in Appendix A additionally serves as a confirmation for policy improvement of regulated segment mixture, which will be elaborated upon later.
@@ III-A System Model
As illustrated in Fig. 1, we primarily consider an IoV scenario consisting of $N$ CAVs, in which each CAV is controlled by an independent SAC-empowered agent, alongside some HDVs. Our MASAC learning encompasses an independent local learning phase and a communication-assisted mixing phase. In the first phase, we use SAC algorithm for each IRL agent $i$, $i\in\{1,2,\cdots,N\}$. That is, agent $i$ senses partial status $s_{t}^{(i)}$ (e.g., the speed and positions of neighboring vehicles) of the IoV environment, and has its local policy $\pi^{(i)}$ approximated by neural networks, and parameterized by $\theta\in\mathbb{R}^{d}$. It collects samples $\langle s_{t}^{(i)},a_{t}^{(i)},r_{t}^{(i)},s_{t+1}^{(i)}\rangle$ in replay buffer $\mathcal{D}^{(i)}$, and randomly samples a mini-batch $\Phi^{(i)}$ for local independent model updates [^1] Hereafter, for simplicity of representation, we omit the superscript $(i)$ under cases where the mentioned procedure applies for any agent.. Subsequently, in the second phase, each agent $i$ interacts with its one-hop neighbors $j\in\Omega_{i}$ within communication range, so as to reduce the behavioral localities of IRL and improve their cooperation efficiency.
@@ III-A1 Local Learning Phase
Algorithmically, in the local learning phase with SAC, parameterized DNNs are used as approximators for policy and soft $Q$-function. Concretely, we alternate optimizing one policy network $\pi$ parameterized by $\theta$ and two soft $Q$ networks parameterized by $\omega_{1}$ and $\omega_{2}$, respectively. Besides, there are also two target soft $Q$ networks parameterized by $\bar{\omega}_{1}$ and $\bar{\omega}_{2}$, obtained as an exponentially moving average of current $Q$ network weights $\omega_{1}$, $\omega_{2}$. Yet, only the minimum $Q$ value of the two soft $Q$-functions is used for the SGD and policy gradient. This setting of two soft $Q$-functions will speed up training while the use of target $Q$ can stabilize the learning [42, 43, 65]. For training $\theta$ and $\omega\in{\omega_{1},\omega_{2}}$, agent randomly samples a batch of transition tuples from the replay buffer $\mathcal{D}$ and performs SGD. The parameters of each soft $Q$ network $\omega_{x},\forall x=1,2$ are updated through minimizing the soft Bellman residual error, that is,
$\vspace{-0.2em}J_{Q}(\omega_{x})=\frac{1}{2}\mathop{\mathbb{E}}\nolimits_{s_{t},a_{t}\sim\mathcal{D}}\left[Q_{\omega_{x}}({s}_{t},{a}_{t})-\hat{Q}(s_{t},a_{t})\right]^{2}\vspace{-0.2em}$ (6)
@@
$J_{\pi}(\theta)=\!\mathop{\mathbb{E}}\limits_{{s}_{t}\sim\mathcal{D}}\!\big{[}\!\mathop{\mathbb{E}}\limits_{{a}_{t}\sim\pi_{\theta}}\!\big{[}\alpha\log\pi_{\theta}({a}_{t}|{s}_{t})\!-\!\mathop{\min}_{\begin{subarray}{c}x\in 1,2\end{subarray}}Q_{\omega_{x}}({s}_{t},{a}_{t})\big{]}\big{]}\vspace{-0.2em}$ (7)
@@
$\vspace{-0.2em}a_{t}=f_{\theta}(s_{t};\delta_{t})=\tanh\big{(}\mu_{\theta}(s_{t})+\delta_{t}\odot\sigma_{\theta}(s_{t})\big{)}\vspace{-0.2em}$ (8)
where $\odot$ represents Hadamard product, $\delta_{t}$ is sampled from $\mathcal{N}(0,\mathbf{I}_{\dim\mathcal{A}})$, the mean $\mu_{\theta}$ and standard $\sigma_{\theta}$ are outputs from the policy network parameterized by $\theta\in\mathbb{R}^{d}$. This reparameterization enables the action $a_{t}$ to be a differentiable function of $\theta$, facilitating gradient descent methods.
@@
$J(\alpha)=\mathop{\mathbb{E}}\limits_{s_{t}\sim\mathcal{D}}\big{[}\mathop{\mathbb{E}}_{{a}_{t}\sim\pi_{\theta}}[-\alpha\log\pi_{\theta}({a}_{t}|{s}_{t})-\alpha\bar{\mathcal{H}}]\big{]}\vspace{-0.2em}$ (9)
where $\bar{\mathcal{H}}$ is an entropy target with default value $-\dim{\mathcal{A}}$. Thus the policy can explore more in regions where the optimal action is uncertain, and remain more deterministic in states with a clear distinction between good and bad actions.
In addition, since two-timescale updates, i.e, less frequent policy updates, usually result in higher quality policy updates. Therefore, we integrate the delayed policy update mechanism employed in TD3 [66], into our framework’s methodology. To this end, the policy, temperature and target $Q$ networks are updated with respect to soft $Q$ network every $e$ iterations.
@@ III-A2 Communication-Assisted Mixing Phase
Subsequent to the phase of local learning, neighboring agents initiate periodic communication to enhance their collaboration in accomplishing complex tasks. The messages transmitted by agents are limited to policy parameters $\theta$ as in VFL [17, 18, 19]. Such an assumption is feasible as soft $Q$-functions have less impact on action selection than the policy in actor-critic algorithm.
Specifically, every $U$ times of policy updates, a communication round begins. Agent $i$ receives the policy parameters $\theta^{(j)}$ from neighboring agents $j\in\Omega_{i}$ via the V2V collaboration channel. Subsequently, agent $i$ could employ various methods to formulate a referential policy $\tilde{\pi}^{(i)}$, which is parameterized by $\tilde{\theta}^{(i)}\!=\!f({\theta}^{(1)},\cdots,{\theta}^{(j)},\cdots)$, that is, the parameters obtained from its neighbors $\forall j\in\Omega_{i}$.
Afterwards, agent $i$ directly mixes DNN parameters $\theta$ of the policy network, consistently with parallel distributed SGD methods as
$\theta_{\text{mix}}^{(i)}=\theta^{(i)}+\zeta(\tilde{\theta}^{(i)}-\theta^{(i)})\vspace{-0.1em}$ (10)
where $\zeta\in[0,1]$ is the mixture metric of DNN parameters. Taking model averaging in [35, 44] as example, $\tilde{\theta}^{(i)}$ is computed as $\tilde{\theta}^{(i)}\!=\!\sum\nolimits_{j\in\Omega_{i}}\!{\theta}^{(j)}$, and $\zeta\!=\!1\!-\!1/|\Omega_{i}|$ is further influenced by the number of neighbors involved. Then, for each agent $i$, $\theta^{(i)}$ should get aligned with mixed policy’s parameters $\theta_{\text{mix}}^{(i)}$.
@@ III-B Problem Formulation
This paper primarily targets the communication assisted mixing phase in MASAC. Instead of simply continuing to follow the idea of direct average of the DNN parameters in FL, an effective parameter mixture method could better leverage the exchanged parameters to yield a superior target referential policy. This approach aims to improve the reinforcement learning policy performance during communication assisted mixing phase, resulting in a consistently higher value with respect to the maximum entropy objective, as described in (1). However, it remains little investigated on the feasible means to mix the exchanged parameters (or their partial segments) and determine the proper mixture metric in (10), though it vitally affects both the communication overhead and learning performance. Therefore, by optimizing the mixture metric $\zeta$, we mainly focus on reducing the communication expenditure while maintain acceptable cumulative rewards, that is,
@@
$s.t.\quad$ $\sum\nolimits_{t}r_{t}(\Theta,\zeta)\geq r^{\text{thre}}$
$\Theta\leftarrow\{\theta_{\text{mix},k}^{(1)},\cdots,\theta_{\text{mix},k}^{(N)}\}$ (11)
$\theta_{\text{mix},k}^{(i)}=\theta_{k}^{(i)}+\zeta(\tilde{\theta}_{k}^{(i)}-\theta_{k}^{(i)}),$ $\forall i\in\{1,\cdots,N\}$
$\tilde{\theta}_{k}^{(i)}=f({\theta}_{k}^{(1)},\cdots,{\theta}_{k}^{(j)},\cdots),$ $\forall k\text{ mod }U=0,j\in\Omega_{i}$
where $r^{\text{thre}}$ denotes the required minimum cumulative rewards, $v$ indicates the transmission bits of policy parameters, and $k$ is the index of policy iterations. Furthermore, $c(v,f)$ denotes the communication expenditure, governed by the utilization function $f$, which represents the method of utilizing the parameter or parameter segments obtained from different neighboring agent $j\in\Omega_{i}$ to construct a referential policy parameterized by $\tilde{\theta}$. The practical implementation of $f$ can be contingent on various factors, i.e., the underlying communications capability of vehicles to simultaneously receive and handle signals and the environmental conditions. After obtaining the referential policy parameter $\tilde{\theta}$, it is worthwhile to resort to a more comprehensive design of $\zeta$ to calibrate the communicating agents and contents as well as regulate the means to mix parameters or parameter segments, so as to provide a guarantee of performance improvement.
@@ IV Mixed Performance Improvement Bound Theorem of FRL Communication under MERL
Given the potential for significant variability in policy performance due to differences in training samples among multiple agents, as well as the staleness of iterations caused by varying computing power of different vehicles, it is critical to recognize that not all referential policies - those amalgamated with the agent’s own policy via the DNN parameter mixture approach detailed in (10) - can contribute positively to agent’s local learning process. More seriously, it may even degrade the learning performance sometimes [44, 46]. Therefore, in order to ensure the policy improvement post-mixture, a robust theoretical analysis method for evaluating the efficacy of the mixed policy distribution $\pi_{\text{mix}}$, approximated by DNN with parameter $\theta_{\text{mix}}$, is essential. This would regulate the selection of only those referential policies $\tilde{\pi}$, parameterized by $\tilde{\theta}$, that are beneficial, into the parameter mixture process.
@@
$\pi_{\text{mix}}(a|s)=(1-\beta)\pi(a|s)+\beta\tilde{\pi}(a|s)$ (12)
@@
$\quad+\alpha\mathop{\mathbb{E}}_{s\sim d_{\pi_{\text{mix}}}}\left[\mathrm{D}_{\mathrm{JS}}^{\beta}(\tilde{\pi}(\cdot|s)\|\pi(\cdot|s))\right]$ (13)
@@
Furthermore, since for $\beta\in[0,1]$, $\mathrm{D}_{\text{JS}}^{\beta}$ is greater than zero [68], only the sign of the first two terms in (1) need to be considered. By applying the re-defined policy advantage under MERL
$\mathbb{A}^{+}_{\pi}(\tilde{\pi}):=\mathop{\mathbb{E}}_{\begin{subarray}{c}s\sim d_{\pi},a\sim\tilde{\pi}\end{subarray}}\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]$ (14)
@@
$\eta(\pi_{\text{mix}})-\eta(\pi)$ $\geq\beta\mathbb{A}^{+}_{\pi}(\tilde{\pi})-C\beta^{2}$ (15)

### v3

@@ I-D Paper Organization
The remainder of this paper is organized as follows. In Section II, we present preliminaries of SAC algorithm and main notations used in this paper. Then, we introduce the system model and formulate the problem in Section III. Afterward, we provide the mixed performance improvement bound theorem of FRL communication under MERL in Section IV, and elaborate on the details of the proposed RSM-MASAC algorithm in Section V. In Section VI, we present the simulation settings and discuss the experimental results. Finally, Section VII concludes this paper and discusses future work.
@@ II Preliminary
$s_{t}^{(i)},a_{t}^{(i)},r^{(i)}_{t}$ Local state, individual action and reward of agent $i$ at time step $t$.
@@
$p,P$ Index of segments and segmentation granularity, $p=1,2,\cdots,P$.
$\Omega_{i}$ Set of one-hop neighbors within the communication range of agent $i$.
$\zeta$ Mixture metric of current policy parameter vector and referential policy parameter vector.
$\alpha$ Temperature parameter of policy entropy.
$\varrho$ Target smoothing coefficient of target $Q$ networks.
$\kappa$ Predefined model replica number.
$U$ Communication interval determined by specified iterations of the local policy.
$\upsilon$ Transmission bits of the policy parameters.
$\psi$   Communication consumption.
@@
$\vspace{-0.2em}\max\eta(\pi)=\mathop{\mathbb{E}}_{s_{0}\sim\rho_{0}}\left[V^{\pi}(s_{0})\right],$ (1)
@@
    $V(s_{t})$ $=\mathbb{E}_{a_{t}\sim\pi}[Q(s_{t},a_{t})]+\alpha H(\pi(\cdot|s_{t})){.}$ (4)
@@
  • In the soft policy improvement step, the goal is to find a policy $\pi_{\text{new}}$ superior to the current policy $\pi_{\text{old}}$, in terms of maximizing (1). Specifically, for each state, SAC updates the policy as
@@
    $=$ $\mathop{\arg\max}_{\pi\in\Pi}\mathbb{E}_{a_{t}\sim\pi}\left[Q^{\pi_{\text{old}}}(s_{t},a_{t})-\alpha\log\pi(a_{t}|s_{t})\right]{,}$
    where $\mathrm{D}_{\mathrm{KL}}(\cdot)$ denotes the KL divergence, while the partition function $Z^{\pi_{\text{old}}}$ normalizes the distribution. The last equality in (5) is due to that $Z^{\pi_{\text{old}}}$ has no contribution to gradient with respect to the new policy, it can thus be ignored. As unveiled in Appendix A, through the update rule of (5), $Q^{\pi_{\text{new}}}(s_{t},a_{t})\geq Q^{\pi_{\text{old}}}(s_{t},a_{t})$ is guaranteed for all $(s_{t},a_{t})\in\mathcal{S}\times\mathcal{A}$. Notably, the proof of soft policy improvement detailed in Appendix A additionally serves as a confirmation for policy improvement of regulated segment mixture, which will be elaborated upon later.
@@ III-A System Model
We primarily consider a system consisting of $N$ agents empowered by the MASAC learning, which encompasses an independent local learning phase and a communication-assisted mixing phase. In the first phase, we use SAC algorithm for each IRL agent $i$, $i\in\{1,2,\cdots,N\}$. That is, agent $i$ senses partial status $s_{t}^{(i)}$ and has its local policy $\pi^{(i)}$ approximated by neural networks, and parameterized by $\theta\in\mathbb{R}^{d}$. It collects samples $\langle s_{t}^{(i)},a_{t}^{(i)},r_{t}^{(i)},s_{t+1}^{(i)}\rangle$ in replay buffer $\mathcal{D}^{(i)}$, and randomly samples a mini-batch $\Phi^{(i)}$ for local independent model updates [^1] Hereafter, for simplicity of representation, we omit the superscript $(i)$ under cases where the mentioned procedure applies for any agent.. Subsequently, in the second phase, each agent $i$ interacts with its one-hop neighbors $j\in\Omega_{i}$ within communication range, so as to reduce the behavioral localities of IRL and improve their cooperation efficiency.
@@ III-A1 Local Learning Phase
Algorithmically, in the local learning phase with SAC, parameterized DNNs are used as approximators for policy and soft $Q$-function. Concretely, we alternate optimizing one network of policy $\pi$ parameterized by $\theta$ and two soft $Q$ networks parameterized by $\omega_{1}$ and $\omega_{2}$, respectively. Besides, there are also two target soft $Q$ networks parameterized by $\bar{\omega}_{1}$ and $\bar{\omega}_{2}$, obtained as an exponentially moving average of current $Q$ network weights $\omega_{1}$, $\omega_{2}$. Yet, only the minimum $Q$ value of the two soft $Q$-functions is used for the SGD and policy gradient. This setting of two soft $Q$-functions will speed up training while the use of target $Q$ can stabilize the learning [37, 38, 62]. For training $\theta$ and $\omega\in{\omega_{1},\omega_{2}}$, agent randomly samples a batch of transition tuples from the replay buffer $\mathcal{D}$ and performs SGD. The parameters of each soft $Q$ network $\omega_{x},\forall x=1,2$ are updated through minimizing the soft Bellman residual error, that is,
$\vspace{-0.2em}J_{Q}(\omega_{x})=\frac{1}{2}\mathop{\mathbb{E}}\nolimits_{s_{t},a_{t}\sim\mathcal{D}}\left[Q_{\omega_{x}}({s}_{t},{a}_{t})-\hat{Q}(s_{t},a_{t})\right]^{2}{,}\vspace{-0.2em}$ (6)
@@
$J_{\pi}(\theta)=\!\mathop{\mathbb{E}}\limits_{{s}_{t}\sim\mathcal{D}}\!\big{[}\!\mathop{\mathbb{E}}\limits_{{a}_{t}\sim\pi_{\theta}}\!\big{[}\alpha\log\pi_{\theta}({a}_{t}|{s}_{t})\!-\!\mathop{\min}_{\begin{subarray}{c}x\in 1,2\end{subarray}}Q_{\omega_{x}}({s}_{t},{a}_{t})\big{]}\big{]}{.}\vspace{-0.2em}$ (7)
@@
$\vspace{-0.2em}a_{t}=f_{\theta}(s_{t};\delta_{t})=\tanh\big{(}\mu_{\theta}(s_{t})+\delta_{t}\odot\sigma_{\theta}(s_{t})\big{)}{,}\vspace{-0.2em}$ (8)
where $\odot$ represents Hadamard product, $\delta_{t}$ is sampled from $\mathcal{N}(0,\mathbf{I}_{\dim\mathcal{A}})$, the mean $\mu_{\theta}$ and standard $\sigma_{\theta}$ are outputs from the policy network parameterized by $\theta$. This reparameterization enables the action $a_{t}$ to be a differentiable function of $\theta$, facilitating gradient descent methods.
@@
$J(\alpha)=\mathop{\mathbb{E}}\limits_{s_{t}\sim\mathcal{D}}\big{[}\mathop{\mathbb{E}}_{{a}_{t}\sim\pi_{\theta}}[-\alpha\log\pi_{\theta}({a}_{t}|{s}_{t})-\alpha\bar{\mathcal{H}}]\big{]}{,}\vspace{-0.2em}$ (9)
where $\bar{\mathcal{H}}$ is an entropy target with default value $-\dim{\mathcal{A}}$. Thus the policy can explore more in regions where the optimal action is uncertain, and remain more deterministic in states with a clear distinction between good and bad actions. Besides, since two-timescale updates, i.e, less frequent policy updates, usually result in higher quality policy updates, we integrate the delayed policy update mechanism employed in TD3 [63] into our framework’s methodology. To this end, the policy, temperature and target $Q$ networks are updated with respect to soft $Q$ network every $e$ iterations.
@@ III-A2 Communication-Assisted Mixing Phase
Subsequent to the phase of local learning, neighboring agents initiate periodic communication to enhance their collaboration in accomplishing complex tasks. The messages transmitted by agents are limited to policy parameters $\theta$ as in many works [19, 20, 24]. Such an assumption is feasible as soft $Q$-functions have less impact on action selection than the policy in actor-critic algorithms.
Specifically, every $U$ times of policy updates, a communication round begins. Agent $i$ receives the policy parameters $\theta^{(j)}$ from neighboring agents $j\in\Omega_{i}$ via the D2D collaboration channel. Subsequently, agent $i$ could employ various methods to formulate a referential policy $\tilde{\pi}^{(i)}$, which is parameterized by $\tilde{\theta}^{(i)}\!=\!f({\theta}^{(1)},\cdots,{\theta}^{(j)},\cdots)$, that is, the parameters obtained from its neighbors $\forall j\in\Omega_{i}$. Afterward, agent $i$ directly mixes DNN parameters of the policy network, consistently with parallel distributed SGD methods as
$\theta_{\text{mix}}^{(i)}=\theta^{(i)}+\zeta(\tilde{\theta}^{(i)}-\theta^{(i)}){,}\vspace{-0.1em}$ (10)
where $\zeta\in[0,1]$ is the mixture metric of DNN parameters. Taking model averaging in [30, 39] as the example, $\tilde{\theta}^{(i)}$ is computed as $\tilde{\theta}^{(i)}\!=\frac{1}{|\Omega_{i}|}\!\sum\nolimits_{j\in\Omega_{i}}\!{\theta}^{(j)}$, and $\zeta\!=\!1\!-\!1/(|\Omega_{i}|+1)$ is further influenced by the number of neighbors involved. Then, for each agent $i$, $\theta^{(i)}$ should get aligned with mixed policy’s parameters $\theta_{\text{mix}}^{(i)}$.
@@ III-B Problem Formulation
This paper primarily targets the communication-assisted mixing phase. Instead of simply continuing to follow the idea of the direct average of the DNN parameters in FL, an effective parameter mixture method could better leverage the exchanged parameters to yield a superior target referential policy. This approach aims to improve RL policy performance during the communication-assisted mixing phase, resulting in a consistently higher value with respect to the maximum entropy objective, as described in (1). However, it remains little investigated on the feasible means to mix the exchanged parameters (or their partial segments) and determine the proper mixture metric in (10), though it vitally affects both the communication overhead and learning performance. Therefore, by optimizing the mixture metric $\zeta$, we mainly focus on reducing the communication expenditure while maintaining acceptable cumulative rewards, that is,
@@
$s.t.\quad$ $\sum\nolimits_{t}r_{t}(\Theta,\zeta)\geq r^{\text{thre}}{,}$
$\Theta\leftarrow\{\theta_{\text{mix},k}^{(1)},\cdots,\theta_{\text{mix},k}^{(N)}\}{,}$ (11)
$\theta_{\text{mix},k}^{(i)}=\theta_{k}^{(i)}+\zeta(\tilde{\theta}_{k}^{(i)}-\theta_{k}^{(i)}){,}$ $\forall i\in\{1,\cdots,N\}{,}$
$\tilde{\theta}_{k}^{(i)}=f({\theta}_{k}^{(1)},\cdots,{\theta}_{k}^{(j)},\cdots){,}$ $\forall k\text{ mod }U=0,j\in\Omega_{i}{,}$
where $r^{\text{thre}}$ denotes the required minimum cumulative rewards, $v$ indicates the transmission bits of policy parameters, and $k$ is the index of policy iterations. Furthermore, $c(v,f)$ denotes the communication expenditure, governed by the utilization function $f$, which represents the method of utilizing the parameter or parameter segments obtained from different neighboring agent $j\in\Omega_{i}$ to construct a referential policy parameterized by $\tilde{\theta}$. The practical implementation of $f$ can be contingent on various factors, i.e., the underlying communications capability of agents to simultaneously receive and handle signals and the environmental conditions. After obtaining the referential policy parameter $\tilde{\theta}$, it is worthwhile to resort to a more comprehensive design of $\zeta$ to calibrate the communicating agents and contents as well as regulate the means to mix parameters or parameter segments, so as to provide a guarantee of performance improvement.
@@ IV Mixed Performance Improvement Bound Theorem of FRL Communication under MERL
Given the potential for significant variability in policy performance due to differences in training samples among multiple agents, as well as the staleness of iterations caused by the varying computing power of different agents, it is critical to recognize that not all referential policies - those amalgamated with the agent’s own policy via the DNN parameter mixture approach detailed in (10) - can contribute positively to agent’s local learning process. More seriously, it may even degrade the learning performance sometimes [39, 41]. Therefore, in order to ensure the policy improvement post-mixture, a robust theoretical analysis method for evaluating the efficacy of the mixed policy distribution $\pi_{\text{mix}}$, approximated by DNN with parameter $\theta_{\text{mix}}$, is essential. This would regulate the selection of only those referential policies $\tilde{\pi}$, parameterized by $\tilde{\theta}$, that are beneficial, in the parameter mixture process.
@@
$\pi_{\text{mix}}(a|s)=(1-\beta)\pi(a|s)+\beta\tilde{\pi}(a|s){,}$ (12)
@@
$\quad+\alpha\mathop{\mathbb{E}}_{s\sim d_{\pi_{\text{mix}}}}\left[\mathrm{D}_{\mathrm{JS}}^{\beta}(\tilde{\pi}(\cdot|s)\|\pi(\cdot|s))\right]{,}$ (13)
@@
Furthermore, since for $\beta\in[0,1]$, $\mathrm{D}_{\text{JS}}^{\beta}$ is greater than zero [65], only the sign of the first two terms in (1) needs to be considered. By applying the re-defined policy advantage under MERL
$\mathbb{A}^{+}_{\pi}(\tilde{\pi}):=\mathop{\mathbb{E}}_{\begin{subarray}{c}s\sim d_{\pi},a\sim\tilde{\pi}\end{subarray}}\left[A_{\pi}(s,a)+\alpha H(\tilde{\pi}(\cdot|s))\right]{,}$ (14)
@@
$\eta(\pi_{\text{mix}})-\eta(\pi)$ $\geq\beta\mathbb{A}^{+}_{\pi}(\tilde{\pi})-C\beta^{2}{,}$ (15)